# The effect of accretion on scalar superradiant instability

### v1

@@ V Summary and Discussion
In Sec. II, we reviewed the scalar superradiance rate, the evolution of the BH-condensate system without accretion, and the evolution of BHs with only accretion. The capture of photons from the accretion disk is included, resulting in a limiting value of the BH spin $a_{*\text{lim}}=0.998$. Without superradiance, the BH mass grows exponentially with time and the spin follows the accretion trajectory to $a_{*\text{lim}}$. The differential evolution equations with all these effects are provided in Eqs. (11), which serves as the analysis benchmark in the rest of this manuscript.
@@
In Sec. III, we kept only the $\{0,1,1\}$ mode and study the influence of accretion on this BH-condensate system. Without accretion, the system has two characteristic time scales, i.e., the superradiance time scale $\tau_{\text{SR}}$ and the GW emission time scale $\tau_{\text{GW}}$. In the presence of accretion, there is an additional accretion time scale $\tau_{\text{acc}}$, given in Eq. (18). Generally, the evolution could be split into the spin-up, spin-down, attractor, and decaying phases. In each phase, the unimportant effects could be ignored and analytical estimates of the important quantities have been obtained as a series of $\alpha$. When $\tau_{\text{acc}}\geq\tau_{\text{SR}}$, the evolution is strongly affected by the accretion, especially in the parameter space with small initial mass coupling or small initial BH spin. In particular, the condensate accumulates mass as a double exponential function, leading to a much stronger GW signal than that without accretion. The lifetime of the mode and the duration of the GW emission are greatly reduced, which is in agreement with the findings of Ref. Brito:2014wla .
@@
Then in Sec. IV, we generalized the analysis of the single-mode case to that with multiple modes. The $\{n,l,l\}$ modes with $n=0,1$ and $0\leq l\leq 7$ are included. The evolution on the Regge plot presents a quite simple and universal pattern. It consists of the parts following along the accretion trajectory or different $\{0,l,l\}$ Regge trajectories, as well as the transitions between these curves, see Fig. 6. Analytical approximations of the start and end of each transition are provided, which agree reasonably well with the numerical results. Same as the single-mode case, the accretion impact strongly on those modes with superradiance time scale $\tau_{\text{SR}}$ equal to or larger than the accretion time scale $\tau_{\text{acc}}$. The accretion enhances the magnitude and reduces the duration of the GW emission from these modes.
@@
In this work, we did not consider the back-reaction of the scalar condensate on the spacetime metric as well as the self-interaction of the scalar field. The self-interaction could generate much richer physics in the evolution, such as the bosonova collapse in Refs. Yoshino:2012kn ; Fukuda:2019ewf ; Omiya:2022mwv ; Omiya:2022gwu ; Unal:2023yxt ; Omiya:2024xlz . Moreover, the coexistence of an accretion disk and a scalar condensate may lead to new observational signatures if their interactions exist. It has been shown that interaction between scalar and photons could modify the cosmic microwave background Blas:2020nbs or result in a birefringent effect of the background light Chen:2019fsq ; Chen:2021lvo .
@@
Finally, we have assumed the dominant mode is always more massive than the subdominant one with the same azimuthal number $l$. Not considering the unnatural cases in which the subdominant mode has a much larger initial mass, this statement is strictly valid only for $l\leq 2$. For $l\geq 3$, there exists the “overtone mixing” phenomenon which says the superradiance rate of the subdominant mode is larger than the dominant mode in some parameter space. Consequently, the subdominant mode is more massive even when its initial mass is not very large. Next, we argue the above approximations for the BH evolution could still give reasonable results. With overtone mixing, both the dominant and subdominant modes should be kept. Since $\omega_{0ll}\approx\omega_{1ll}$, the second piece on the right side of Eq. (11b) is roughly $-l\,\omega_{0ll}^{-1}\,(\dot{E}_{\text{SR}}^{(0ll)}+\dot{E}_{\text{SR}}^{(1ll)})$. The Eq. (11c) for these two modes could also be added. The obtained differential equations are almost the same as Eqs. (11), except the $M_{\text{s}}^{(0ll)}$ and $\dot{E}_{\text{SR}}^{(0ll)}$ are replaced by $M_{\text{s}}^{(0ll)}+M_{\text{s}}^{(1ll)}$ and $\dot{E}_{\text{SR}}^{(0ll)}+\dot{E}_{\text{SR}}^{(1ll)}$, respectively. With these two replacements, the later analysis is unchanged, which results in the same evolution of the BH on the Regge plot. For the $l=4$ stage, the overtone mixing introduces an additional percentage error of roughly $10\%$ compared to the numerical calculation. Nonetheless, since the dominant and subdominant modes have very different GW emission rates, the GW emission flux will be largely modified with overtone mixing. We leave the related analysis in a future publication.
@@ Appendix A Some useful formulae
For particles in the innermost stable circular orbit, the energy per mass $E_{\mathrm{ms}}^{\dagger}$ and angular momentum per mass $J_{\mathrm{ms}}^{\dagger}$ are given by Bardeen:1970zz ; Bardeen:1972fi
$E_{\mathrm{ms}}^{\dagger}$ $=\sqrt{1-\frac{2}{3r_{\mathrm{ms}}/M}},$ (78a)
$J_{\mathrm{ms}}^{\dagger}$ $=\frac{2M}{3\sqrt{3}}\left(1+2\sqrt{3r_{\mathrm{ms}}/M-2}\right),$ (78b)
@@
$r_{\mathrm{ms}}/M$ $=3+z_{2}-\sqrt{(3-z_{1})(3+z_{1}+2z_{2})},$ (79)
[…]
$z_{2}$ $=\sqrt{3a_{*}^{2}+z_{1}^{2}}.$ (81)
@@
$\dot{E}_{\mathrm{GW}}^{(0ll)}={C}_{\mathrm{GW}}^{(0ll)}\left(\frac{{M}_{\mathrm{s}}^{(0ll)}}{M}\right)^{2}\left(\frac{M\mu}{l}\right)^{4l+10}.$ (82)
@@
${C}_{\mathrm{GW}}^{(011)}$ $\approx 2.49\times 10^{-2},$ (83a)
${C}_{\mathrm{GW}}^{(022)}$ $\approx 7.29\times 10^{-2},$ (83b)
${C}_{\mathrm{GW}}^{(033)}$ $\approx 8.97\times 10^{-2},$ (83c)
${C}_{\mathrm{GW}}^{(044)}$ $\approx 4.29,$ (83d)
${C}_{\mathrm{GW}}^{(055)}$ $\approx 2.24\times 10^{2},$ (83e)
${C}_{\mathrm{GW}}^{(066)}$ $\approx 6.45\times 10^{3},$ (83f)
${C}_{\mathrm{GW}}^{(077)}$ $\approx 1.46\times 10^{5}.$ (83g)
@@
$\dot{E}_{\mathrm{beat}}^{(l)}={C}_{\mathrm{beat}}^{(l)}\frac{\left({M}_{\mathrm{s}}^{(0ll)}\right)^{3/2}\left({M}_{\mathrm{s}}^{(1ll)}\right)^{1/2}}{M^{2}}\left(\frac{M\mu}{l}\right)^{4l+10},$ (84)
@@
${C}_{\mathrm{beat}}^{(1)}$ $\approx 5.89\times 10^{-2},$ (85a)
${C}_{\mathrm{beat}}^{(2)}$ $\approx 2.26\times 10^{-1},$ (85b)
${C}_{\mathrm{beat}}^{(3)}$ $\approx 3.33\times 10^{-1},$ (85c)
${C}_{\mathrm{beat}}^{(4)}$ $\approx 1.82\times 10^{1},$ (85d)
${C}_{\mathrm{beat}}^{(5)}$ $\approx 1.06\times 10^{3},$ (85e)
${C}_{\mathrm{beat}}^{(6)}$ $\approx 3.11\times 10^{4},$ (85f)
${C}_{\mathrm{beat}}^{(7)}$ $\approx 8.11\times 10^{5}.$ (85g)
@@
$\frac{{\tau}_{\mathrm{life}}^{(1ll)}}{M}={C}_{\mathrm{life}}^{(1ll)}\left(\frac{M\mu}{l}\right)^{-4l-8}.$ (86)
@@
${C}_{\mathrm{life}}^{(111)}$ $\approx 1.63\times 10^{2},$ (87a)
${C}_{\mathrm{life}}^{(122)}$ $\approx 3.88,$ (87b)
${C}_{\mathrm{life}}^{(133)}$ $\approx 4.29\times 10^{-1},$ (87c)
${C}_{\mathrm{life}}^{(144)}$ $\approx 8.26\times 10^{-2},$ (87d)
${C}_{\mathrm{life}}^{(155)}$ $\approx 2.10\times 10^{-2},$ (87e)
${C}_{\mathrm{life}}^{(166)}$ $\approx 6.27\times 10^{-3},$ (87f)
${C}_{\mathrm{life}}^{(177)}$ $\approx 2.09\times 10^{-3}.$ (87g)
@@
$M^{(0ll)}_{\text{s}}=M^{(0ll)}_{\text{s0}}\exp\left\{\alpha_{0}^{(4l+4)}\mu\tau_{\text{acc}}\left[g_{0ll}\left(\frac{M}{M_{0}}\right)-g_{0ll}\left(1\right)\right]\right\},\hskip 14.22636pt\text{for }1\leq\frac{M}{M_{0}}\leq{\sqrt{6}\,k}.$ (88)

### v2

@@ 5 Summary and Discussion
In Sec. 2, we reviewed the scalar superradiance rate, the evolution of the BH-condensate system without accretion, and the evolution of BHs with only accretion. The capture of photons from the accretion disk is included, resulting in a limiting value of the BH spin $a_{*\text{lim}}=0.998$. Without superradiance, the BH mass grows exponentially with time and the spin follows the accretion trajectory to $a_{*\text{lim}}$. The differential evolution equations with all these effects are provided in Eqs. (2.11), which serves as the analysis benchmark in the rest of this manuscript.
@@
In Sec. 3, we kept only the $\{0,1,1\}$ mode and study the influence of accretion on this BH-condensate system. Without accretion, the system has two characteristic time scales, i.e., the superradiance time scale $\tau_{\text{SR}}$ and the GW emission time scale $\tau_{\text{GW}}$. In the presence of accretion, there is an additional accretion time scale $\tau_{\text{acc}}$, given in Eq. (2.18). Generally, the evolution could be split into the spin-up, spin-down, attractor, and decaying phases. In each phase, the unimportant effects could be ignored and analytical estimates of the important quantities have been obtained as a series of $\alpha$. When $\tau_{\text{acc}}\geq\tau_{\text{SR}}$, the evolution is strongly affected by the accretion, especially in the parameter space with small initial mass coupling or small initial BH spin. In particular, the condensate accumulates mass as a double exponential function, leading to a much stronger GW signal than that without accretion. The lifetime of the mode and the duration of the GW emission are greatly reduced, which is in agreement with the findings of Refs. [22, 27].
@@
Then in Sec. 4, we generalized the analysis of the single-mode case to that with multiple modes. The $\{n,l,l\}$ modes with $n=0,1$ and $0\leq l\leq 7$ are included. The evolution on the Regge plot presents a quite simple and universal pattern. It consists of the parts following along the accretion trajectory or different $\{0,l,l\}$ Regge trajectories, as well as the transitions between these curves, see Fig. 6. Analytical approximations of the start and end of each transition are provided, which agree reasonably well with the numerical results. Same as the single-mode case, the accretion impact strongly on those modes with superradiance time scale $\tau_{\text{SR}}$ equal to or larger than the accretion time scale $\tau_{\text{acc}}$. The accretion enhances the magnitude and reduces the duration of the GW emission from these modes.
@@
In this work, we did not consider the back-reaction of the scalar condensate on the spacetime metric as well as the self-interaction of the scalar field. The self-interaction could generate much richer physics in the evolution, such as the bosonova collapse in Refs. [38, 23, 39, 40, 26, 41]. Moreover, the coexistence of an accretion disk and a scalar condensate may lead to new observational signatures if their interactions exist. It has been shown that interaction between scalar and photons could modify the cosmic microwave background [7] or result in a birefringent effect of the background light [5, 6].
@@
Finally, we have assumed the dominant mode is always more massive than the subdominant one with the same azimuthal number $l$. Not considering the unnatural cases in which the subdominant mode has a much larger initial mass, this statement is strictly valid only for $l\leq 2$. For $l\geq 3$, there exists the “overtone mixing” phenomenon which says the superradiance rate of the subdominant mode is larger than the dominant mode in some parameter space. Consequently, the subdominant mode is more massive even when its initial mass is not very large. Next, we argue the above approximations for the BH evolution could still give reasonable results. With overtone mixing, both the dominant and subdominant modes should be kept. Since $\omega_{0ll}\approx\omega_{1ll}$, the second piece on the right side of Eq. (2.11b) is roughly $l\,\omega_{0ll}^{-1}\,(\dot{E}_{\text{SR}}^{(0ll)}+\dot{E}_{\text{SR}}^{(1ll)})$. The Eq. (2.11c) for these two modes could also be added. The obtained differential equations are almost the same as Eqs. (2.11), except the $M_{\text{s}}^{(0ll)}$ and $\dot{E}_{\text{SR}}^{(0ll)}$ are replaced by $M_{\text{s}}^{(0ll)}+M_{\text{s}}^{(1ll)}$ and $\dot{E}_{\text{SR}}^{(0ll)}+\dot{E}_{\text{SR}}^{(1ll)}$, respectively. With these two replacements, the later analysis is unchanged, which results in the same evolution of the BH on the Regge plot. For the $l=4$ stage, the overtone mixing introduces an additional percentage error of roughly $10\%$ compared to the numerical calculation. Nonetheless, since the dominant and subdominant modes have very different GW emission rates, the GW emission flux will be largely modified with overtone mixing. We leave the related analysis in a future publication.
@@ Appendix A Some useful formulae
For particles in the innermost stable circular orbit, the energy per mass $E_{\mathrm{ms}}^{\dagger}$ and angular momentum per mass $J_{\mathrm{ms}}^{\dagger}$ are given by [42, 43]
$E_{\mathrm{ms}}^{\dagger}$ $=\sqrt{1-\frac{2}{3r_{\mathrm{ms}}/M}},$ (A.1a)
$J_{\mathrm{ms}}^{\dagger}$ $=\frac{2M}{3\sqrt{3}}\left(1+2\sqrt{3r_{\mathrm{ms}}/M-2}\right),$ (A.1b)
@@
$r_{\mathrm{ms}}/M$ $=3+z_{2}-\sqrt{(3-z_{1})(3+z_{1}+2z_{2})},$ (A.2)
$z_{1}$ $=1+\sqrt[3]{1-a_{*}^{2}}\left(\sqrt[3]{1+a_{*}}+\sqrt[3]{1-a_{*}}\right),$ (A.3)
$z_{2}$ $=\sqrt{3a_{*}^{2}+z_{1}^{2}}.$ (A.4)
@@
$\dot{E}_{\mathrm{GW}}^{(0ll)}={C}_{\mathrm{GW}}^{(0ll)}\left(\frac{{M}_{\mathrm{s}}^{(0ll)}}{M}\right)^{2}\left(\frac{M\mu}{l}\right)^{4l+10}.$ (A.5)
@@
${C}_{\mathrm{GW}}^{(011)}$ $\approx 2.49\times 10^{-2},$ (A.6a)
${C}_{\mathrm{GW}}^{(022)}$ $\approx 7.29\times 10^{-2},$ (A.6b)
${C}_{\mathrm{GW}}^{(033)}$ $\approx 8.97\times 10^{-2},$ (A.6c)
${C}_{\mathrm{GW}}^{(044)}$ $\approx 4.29,$ (A.6d)
${C}_{\mathrm{GW}}^{(055)}$ $\approx 2.24\times 10^{2},$ (A.6e)
${C}_{\mathrm{GW}}^{(066)}$ $\approx 6.45\times 10^{3},$ (A.6f)
${C}_{\mathrm{GW}}^{(077)}$ $\approx 1.46\times 10^{5}.$ (A.6g)
@@
$\dot{E}_{\mathrm{beat}}^{(l)}={C}_{\mathrm{beat}}^{(l)}\frac{\left({M}_{\mathrm{s}}^{(0ll)}\right)^{3/2}\left({M}_{\mathrm{s}}^{(1ll)}\right)^{1/2}}{M^{2}}\left(\frac{M\mu}{l}\right)^{4l+10},$ (A.7)
@@
${C}_{\mathrm{beat}}^{(1)}$ $\approx 5.89\times 10^{-2},$ (A.8a)
${C}_{\mathrm{beat}}^{(2)}$ $\approx 2.26\times 10^{-1},$ (A.8b)
${C}_{\mathrm{beat}}^{(3)}$ $\approx 3.33\times 10^{-1},$ (A.8c)
${C}_{\mathrm{beat}}^{(4)}$ $\approx 1.82\times 10^{1},$ (A.8d)
${C}_{\mathrm{beat}}^{(5)}$ $\approx 1.06\times 10^{3},$ (A.8e)
${C}_{\mathrm{beat}}^{(6)}$ $\approx 3.11\times 10^{4},$ (A.8f)
${C}_{\mathrm{beat}}^{(7)}$ $\approx 8.11\times 10^{5}.$ (A.8g)
@@
$\frac{{\tau}_{\mathrm{life}}^{(1ll)}}{M}={C}_{\mathrm{life}}^{(1ll)}\left(\frac{M\mu}{l}\right)^{-4l-8}.$ (A.9)
@@
${C}_{\mathrm{life}}^{(111)}$ $\approx 1.63\times 10^{2},$ (A.10a)
${C}_{\mathrm{life}}^{(122)}$ $\approx 3.88,$ (A.10b)
${C}_{\mathrm{life}}^{(133)}$ $\approx 4.29\times 10^{-1},$ (A.10c)
${C}_{\mathrm{life}}^{(144)}$ $\approx 8.26\times 10^{-2},$ (A.10d)
${C}_{\mathrm{life}}^{(155)}$ $\approx 2.10\times 10^{-2},$ (A.10e)
${C}_{\mathrm{life}}^{(166)}$ $\approx 6.27\times 10^{-3},$ (A.10f)
${C}_{\mathrm{life}}^{(177)}$ $\approx 2.09\times 10^{-3}.$ (A.10g)
@@
$M^{(0ll)}_{\text{s}}=M^{(0ll)}_{\text{s0}}\exp\left\{\alpha_{0}^{(4l+4)}\mu\tau_{\text{acc}}\left[g_{0ll}\left(\frac{M}{M_{0}}\right)-g_{0ll}\left(1\right)\right]\right\},\hskip 14.22636pt\text{for }1\leq\frac{M}{M_{0}}\leq{\sqrt{6}\,k}.$ (A.11)